# Improved bounds for randomized Schatten norm estimation of numerically low-rank matrices

### v2

@@ Abstract
In this work, we analyze the variance of a stochastic estimator for computing Schatten norms of matrices. The estimator extracts information from a single sketch of the matrix, that is, the product of the matrix with a few standard Gaussian random vectors. While this estimator has been proposed and used in the literature before, the existing variance bounds are often pessimistic. Our work provides a sharper upper bound on the variance and we also give estimates of the variance that work well for numerically low-rank matrices. Our theoretical findings are supported by numerical experiments, demonstrating that the new bounds are significantly tighter than the existing ones.
@@ 1 Introduction
A variety of variance reduction techniques has been developed, for instance the combination of Hutchinson trace estimator with randomized low-rank approximation; see, e.g., [3, 9, 11]. Algorithms specifically tailored to Schatten-$p$ norm estimation have been discussed in [2, 4]. When computing quadratic forms involving $(A^{T}A)^{p}$, the matrix $A$ needs to be accessed at least $p$ times. This number can be slightly reduced if, instead of computing exact quadratic forms, these are approximated with Chebyshev polynomials as in [2, 4].
If, instead, we are only allowed to access the matrix $A$ once, we can use the Schatten-$2p$ norm estimator proposed by Kong and Valiant [6], an unbiased estimator for $\|A\|_{2p}^{2p}$. The problem considered in [6], restricted to the Gaussian setting, is to extract information on the eigenvalues $\lambda_{1},\ldots,\lambda_{n}$ of a covariance matrix $B=A^{T}A$ given access to some samples drawn from the multivariate Gaussian distribution corresponding to $B$. This is approached via the methods of moments, which requires estimating the $p$-th moment of the spectrum of $B$ defined as $\sum_{i=1}^{n}\lambda_{i}^{p}$. In the language of randomized numerical linear algebra, Kong and Valiant’s algorithm [6, Algorithm 1] estimates Schatten-$2p$ norms of a matrix $A$ in the so-called streaming setting. The samples drawn from the distributions correspond to a sketch of $A$, that is, to a matrix $Y:=A\Omega$ where $\Omega\in\mathbb{R}^{n\times k}$ consists of i.i.d. standard Gaussian entries, for some $k\ll n$. Throughout the paper, we denote the estimator for $\|A\|_{2p}^{2p}$ from [6] as $\hat{\theta}_{2p}$ and summarize its construction procedure in Algorithm 1. In Algorithm 1, line 2 uses MATLAB function notation triu($\cdot$,1) to indicate the strictly upper triangular part of a matrix.
@@
0:  Sketch $Y=A\Omega\in\mathbb{C}^{m\times k}$ and integer $p$
@@
2:  Define $T=\mathrm{triu}(Z,1)$ the strictly upper triangular part of the matrix $Z$.
@@
It was proved in [6, Proposition 4] that the variance of $\hat{\theta}_{2p}$ with a Gaussian sketching matrix $\Omega$ satisfies the following bound:
@@
However, this bound usually severely overestimates the true variance of $\hat{\theta}_{2p}$. As an illustrative example, Figure 1 plots the variance for $\hat{\theta}_{6}$ (estimator of the $6$-th power of the Schatten-$6$ norm) on a $10\times 10$ matrix with singular values $1,1/2,1/3,\ldots,1/10$, together with the bound (2). The bound overestimates the variance by several orders of magnitude.
Figure 1: Comparison between true variance of $\hat{\theta}_{6}$ and the variance bound (2) for a $10\times 10$ matrix $A$.
We will show in Theorem 3 that, with a Gaussian sketching matrix $\Omega$, the variance of $\hat{\theta}_{2p}$ has first-order expansion
@@
Remark 1.
We remark that the variance of $\hat{\theta}_{2p}$ is usually worse than the variance of estimators based on the Hutchinson trace estimator, though both are unbiased. Specifically, since the variance of the vanilla Hutchinson trace estimator [5] for a matrix $B$ with Gaussian random vectors is $2\|B\|_{F}^{2}$, applying it to (1) gives an unbiased estimator of $\|A\|_{2p}^{2p}$ with variance
$\frac{2p}{k}\|A^{2p}\|_{F}^{2}=\frac{2p}{k}\|A\|_{4p}^{4p},$
which is smaller than (3). We remark that, in order to use the Hutchinson trace estimator, the matrix $A$ needs to be accessed more than once (to apply powers of the matrix to a vector). When the only available information about the matrix $A$ is its sketch $Y$, to the best of the authors’ knowledge, the estimator $\hat{\theta}_{2p}$ proposed in [6] is the only method to estimate the Schatten-$2p$ norms of $A$.
Contributions.
In this paper, we focus on the streaming setting and we are interested in better understanding the variance of $\hat{\theta}_{2p}$ and its behavior on matrices with strongly decaying singular values (i.e., numerically low-rank matrices). We derive an improved bound on the variance of $\hat{\theta}_{2p}$ in the case of a Gaussian sketch. Our bounds leverage the rotational invariance of the Gaussian distribution to simplify the analysis of the estimator, and the results we obtain only depend on the singular values of $A$, highlighting a better behavior in the case of rapid singular value decay. Numerical results show that this bound is relatively tight for matrices with rapid singular value decay using moderately large sketch size $k$. Moreover, in all cases our variance bound is more precise than the previous bound (2). We provide a first-order and second-order expansion of the variance of $\hat{\theta}_{2p}$ in $1/k$. The experiments verify that our first-order and second-order estimates capture the behavior of $\hat{\theta}_{2p}$ for matrices that exhibit a strong singular value decay, for moderate values of $k$.
@@
Section 2 reviews the building blocks of $\hat{\theta}_{2p}$. Section 3 presents our main results on the variance of $\hat{\theta}_{2p}$. In Section 4, we illustrate the performance of our bounds and estimates on a variety of examples with/without singular value decay. Our conclusions are summarized in Section 5.
@@ 2 Notations and review of estimator $\hat{\theta}_{2p}$
$Z_{\tau}=\prod_{\ell=1}^{k}Z_{i_{\ell},i_{\ell+1}},$
with the convention that $i_{k+1}\equiv i_{1}$ for simpler notation. With this notation, we are ready to state one key idea towards the construction of $\hat{\theta}_{2p}$.
Proposition 2 ([7] and [6, Fact 2]).
For any $p$-cycle $\tau=(i_{1},\ldots,i_{p})$ with $i_{\ell}\in[k]$, a real matrix $A\in\mathbb{R}^{m\times n}$, and an $n\times k$ random matrix $\Omega=(\omega_{j,\ell})$ with i.i.d. entries with mean 00 and variance $1$,
@@
Setting $Y=A\Omega$, Proposition 2 says that $(Y^{T}Y)_{\tau}$ is an unbiased estimate of $\|A\|_{2p}^{2p}$ for any (fixed) $p$-cycle $\tau$. A natural approach to reduce the variance of this estimator is to average over all possible $p$-cycles. However, this strategy poses significant computational challenges, as naively iterating over all $\frac{k!}{(k-p)!}$ possible cycles becomes computationally infeasible, even for moderately large values of $k$. A surprising linear algebra fact is that the average over all increasing $p$-cycles, that is, $p$-cycles such that $i_{1}<i_{1}<\ldots<i_{p}$, can be efficiently computed by simple linear algebra operations, and this defines $\hat{\theta}_{2p}$ presented in Algorithm 1:
$\hat{\theta}_{2p}=\binom{k}{p}^{-1}\sum_{1\leq i_{1}<\cdots<i_{p}\leq k}(\Omega^{T}A^{T}A\Omega)_{\tau}=\binom{k}{p}^{-1}\mathrm{trace}(T^{p-1}Z),$ (4)
where $Z$ and $T$ are defined as in Algorithm 1. The proof for last identity in (4) can be found in [6, Lemma 1]. We note that Proposition 2 together with the relation (4) guarantee the unbiasedness of $\hat{\theta}_{2p}$.
@@ 3.1 Main results
Our first result provides the first order expansion in $1/k$ for the variance of $\hat{\theta}_{2p}$ with $k$ being the sketch size (Theorem 3).
@@
$\mathrm{Var}(\hat{\theta}_{2p})=\frac{2p^{2}\|A\|_{4p}^{4p}}{k}+\mathcal{O}\left(\frac{1}{k^{2}}\right).$
Moreover, the constant in $\mathcal{O}(\cdot)$ is a symmetric function of the singular values of $A$.
The next theorem (Theorem 4) provides an upper bound for the variance of a scaling of $\hat{\theta}_{2p}$. The bound is a function of $k$, $p$, and Schatten norms of $A$, highlighting lower variance for numerically low-rank matrices.
@@
Let $\hat{\Theta}_{2p}\coloneqq\binom{k}{p}\hat{\theta}_{2p}$. Then the variance of $\hat{\Theta}_{2p}$ is bounded by
@@
$\mathrm{Var}(\hat{\theta}_{2p})\leq\frac{2p^{2}\|A\|_{4p}^{4p}}{k}+\frac{p^{2}(p-1)^{2}}{k^{2}}\left(\|A\|_{4p}^{4p}+\frac{3}{2}\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}-\frac{1}{2}\|A\|_{2p}^{4p}\right)+\mathcal{O}\left(\frac{1}{k^{3}}\right).$

### v3

@@ Abstract
In this work, we analyze the variance of a stochastic estimator for computing Schatten norms of matrices. The estimator extracts information from a single sketch of the matrix, that is, the product of the matrix with a few standard Gaussian random vectors. While this estimator has been proposed and used in the literature before, the existing variance bounds are often pessimistic. Our work provides a new upper bound and estimates of the variance of this estimator. These theoretical findings are supported by numerical experiments, demonstrating that the new bounds are significantly tighter than the existing ones in the case of numerically low-rank matrices.
@@ 1 Introduction
A variety of variance reduction techniques has been developed, for instance the combination of Hutchinson trace estimator with randomized low-rank approximation; see, e.g., [6, 14, 16]. Algorithms specifically tailored to Schatten-$p$ norm estimation have been discussed in [5, 7]. When computing quadratic forms involving $(A^{T}A)^{p}$, at least $p$ matrix-vector products (matvecs) with $A$ are required and must be applied sequentially. This number can be slightly reduced if the quadratic forms themselves are approximated with Chebyshev polynomials as in [5, 7]. Another line of works [2, 3] assumes that the rows of $A$ are available one at a time – this is known as the row-order model. Under this setting, a Schatten-$2p$ norm estimator can be computed after one or multiple passes over the matrix $A$.
The sketching model assumes access to a sketch of the matrix $A$, that is, a matrix $Y:=A\Omega$ for some random matrix $\Omega\in\mathbb{R}^{n\times k}$, but not necessarily assumes the access to the matvec with $A$ and to the entries of $A$. This setting applies, for instance, to a problem considered in [10]: approximating the set of the eigenvalues $\sigma_{1}^{2},\ldots,\sigma_{n}^{2}$ of an unknown covariance matrix $\Sigma=A^{T}A$ given access to $k\ll n$ samples drawn from a multivariate distribution corresponding to $\Sigma$. This problem can be reduced, via the methods of moments [1, 11, 17], to estimate the $p$-th moment of the spectrum of the covariance matrix $\Sigma$, defined as $\sum_{i=1}^{n}\sigma_{i}^{2p}=\|A\|_{2p}^{2p}$. In our work, we focus on the case where the $k$ samples are drawn from a multivariate Gaussian distribution. Such samples correspond to a Gaussian sketch $Y:=A\Omega$ where $\Omega\in\mathbb{R}^{n\times k}$ consists of i.i.d. standard Gaussian entries. In the language of randomized numerical linear algebra, the problem from [10] can be reformulated as estimating the Schatten-$2p$ norms of a matrix $A$ given a Gaussian sketch of $A$. In case the matrix $A$ is known, but very large, the sketching model can also be advantageous since it allows for parallelization of the matrix-vector multiplications with $A$.
The trace estimation techniques mentioned above are not applicable under the sketching model, so other strategies need to be employed. Li, Nguyen, and Woodruff [12] showed that it is possible to extract an unbiased estimate of $\|A\|_{2p}^{2p}$ from a sketch of $A$. Kong and Valiant [10] proposed a computationally efficient way to compute a similar unbiased estimator, which is the focus of our work. In Algorithm 1 we summarize the procedure for constructing the estimator for $\|A\|_{2p}^{2p}$ proposed in [10], using a Gaussian sketch. The estimator is denoted by $\hat{\theta}_{2p}$. In Algorithm 1, line 2 uses the MATLAB function notation triu($\cdot$,1) to indicate the strictly upper triangular part of a matrix. Note that we assume $p\leq k$, otherwise $T^{p-1}=0$ and the estimator would be identically zero.
@@
0:  Gaussian sketch $Y=A\Omega\in\mathbb{C}^{m\times k}$ and positive integer $p\leq k$
@@
2:  Define $T=\texttt{triu}(Z,1)$ as the strictly upper triangular part of the matrix $Z$.
@@
It was proved in [10, Proposition 4] that the variance of $\hat{\theta}_{2p}$ generated by Algorithm 1 satisfies the following bound:
@@
Lower bounds on Schatten norm estimation [12] imply that the dependence on $n$ and $k$ featured in (2) cannot be improved in the worst case. However, the bound (2) severely overestimates the true variance of $\hat{\theta}_{2p}$ for many matrices in practice. As an illustrative example, Figure 1 plots the variance for $\hat{\theta}_{6}$ (estimator of the $6$-th power of the Schatten-$6$ norm) on a $10\times 10$ matrix with singular values $1,1/2,1/3,\ldots,1/10$, together with the bound (2). Note that the bound (2) overestimates the variance by several orders of magnitude.
Figure 1: Comparison between the true variance of $\hat{\theta}_{6}$, the variance bound (2), and our bound from Theorem 3 for a $10\times 10$ matrix $A$. Note that the $y$-axis has been “cut” for improved readability.
Contributions.
In this paper, we are interested in better understanding the variance of $\hat{\theta}_{2p}$ and its behavior on matrices with strongly decaying singular values (i.e., numerically low-rank matrices). We derive a bound on the variance of $\hat{\theta}_{2p}$ that improves (2) by a large margin for all the matrices we have considered; see Figure 1 above for an illustrative example. While the large constants in the bound (2) were due, in part, to the fact that the analysis was done in a more general case, in which the entries of $\Omega$ could be any centered random variable with unit variance and bounded fourth moment, our bounds instead leverage the rotational invariance of the Gaussian distribution to simplify the analysis of the estimator. The results we obtain do not depend explicitly on the size of the matrix, and they only depend on the singular values of $A$. Numerical results show that our bound is relatively tight for matrices with rapid singular value decay using moderately large sketch size $k$. Moreover, in all the numerical experiments our variance bound is much more precise than the previous bound (2). We provide a first-order and second-order expansion of the variance of $\hat{\theta}_{2p}$ in $1/k$. In particular, we will show in Theorem 2 that, with a Gaussian sketching matrix $\Omega$, the variance of $\hat{\theta}_{2p}$ has first-order expansion
@@
The experiments verify that our first-order and second-order estimates capture the behavior of $\hat{\theta}_{2p}$ for matrices that exhibit a strong singular value decay, for moderate values of $k$.
@@
Section 2 reviews the building blocks of $\hat{\theta}_{2p}$. Section 3 presents our main results on the variance of $\hat{\theta}_{2p}$. In Section 4, we illustrate the performance of our bounds and estimates on a variety of examples with or without singular value decay. Our conclusions are summarized in Section 5.
@@ 2 Notations and review of estimator $\hat{\theta}_{2p}$
$Z_{\tau}=\prod_{\ell=1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p}}Z_{i_{\ell},i_{\ell+1}},$
with the convention that $i_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p}+1}\equiv i_{1}$ for simpler notation. With this notation, we are ready to state one key idea towards the construction of $\hat{\theta}_{2p}$.
Lemma 1 ([12] and [10, Fact 2]).
For any $p$-cycle $\tau=(i_{1},\ldots,i_{p})$ with $i_{\ell}\in[k]$, a real matrix $A\in\mathbb{R}^{m\times n}$, and an $n\times k$ random matrix $\Omega=(\omega_{j,\ell})$ where $\omega_{j,\ell}$ are i.i.d. entries with mean 00 and variance $1$,
@@
Proof.
By substituting the matrix $A$ with its singular value decomposition and leveraging the rotational invariance of the standard Gaussian distribution, we can assume without loss of generality that $A$ is diagonal and thus $A^{T}A=\operatorname{\mathrm{diag}}(\sigma_{1}^{2},\ldots,\sigma_{n}^{2})$. Let $\delta=(\delta_{1},\ldots,\delta_{p})\in[n]^{p}$. Then
$\left(\Omega^{T}A^{T}A\Omega\right)_{\tau}=\prod_{h=1}^{p}\left(\Omega^{T}A^{T}A\Omega\right)_{i_{h},i_{h+1}}=\prod_{h=1}^{p}\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,i_{h}}\omega_{t,i_{h+1}}\right)=\sum_{\delta\in[n]^{p}}\prod_{h=1}^{p}\sigma_{\delta_{h}}^{2}\omega_{\delta_{h},i_{h}}\omega_{\delta_{h},i_{h+1}}.$
Since $\omega_{i,j}$’s are i.i.d. random variables with mean 00 and variance $1$, the expectation of the product $\prod_{h=1}^{p}\sigma_{\delta_{h}}^{2}\omega_{\delta_{h},i_{h}}\omega_{\delta_{h},i_{h+1}}$ vanishes unless every entry of $\Omega$ appears an even number of times in the product. Since $i_{1},\ldots,i_{p}$ are distinct, each non-zero term must have each $\omega_{\bullet,i_{h}}$ appear exactly twice and thus $\delta_{h}=\delta_{h+1}$. Therefore, the expectation of $\left(\Omega^{T}A^{T}A\Omega\right)_{\tau}$ becomes
$\mathbb{E}[(\Omega^{T}A^{T}A\Omega)_{\tau}]=\sum_{\delta_{1}=1}^{n}\prod_{h=1}^{p}\left(\mathbb{E}[\omega_{\delta_{1},i_{h}}^{2}]\sigma_{\delta_{1}}^{2}\right)=\sum_{\delta_{1}=1}^{n}\sigma_{\delta_{1}}^{2p}=\operatorname{\mathrm{tr}}((A^{T}A)^{p}).\qed$
Setting $Y=A\Omega$, Lemma 1 says that $(Y^{T}Y)_{\tau}$ is an unbiased estimate of $\|A\|_{2p}^{2p}$ for any (fixed) $p$-cycle $\tau$. A natural approach to reduce the variance of this estimator is to average over all possible $p$-cycles. However, this strategy poses significant computational challenges, as naively iterating over all $\frac{k!}{(k-p)!}$ possible cycles becomes computationally infeasible, even for moderately large values of $k$. A surprising linear algebra fact is that the average over all increasing $p$-cycles, that is, $p$-cycles such that $i_{1}<i_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}}<\ldots<i_{p}$, can be efficiently computed by simple linear algebra operations [10], and this defines $\hat{\theta}_{2p}$ presented in Algorithm 1:
$\binom{k}{p}^{-1}\sum_{1\leq i_{1}<\cdots<i_{p}\leq k}(\Omega^{T}A^{T}A\Omega)_{\tau}=\binom{k}{p}^{-1}\mathrm{trace}(T^{p-1}Z){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\hat{\theta}_{2p}},$ (4)
where $Z$ and $T$ are defined as in Algorithm 1. The proof for the first identity in (4) can be found in [10, Lemma 1]. We note that Lemma 1 together with the relation (4) guarantee the unbiasedness of $\hat{\theta}_{2p}$.
@@ 3.1 Main results
Our first result provides the first order expansion in $1/k$ for the variance of $\hat{\theta}_{2p}$ where $k$ is the sketch size.
@@
$\mathrm{Var}(\hat{\theta}_{2p})=\frac{2p^{2}\|A\|_{4p}^{4p}}{k}+\mathcal{O}\left(\frac{1}{k^{2}}\right),$
where the constant in $\mathcal{O}(\cdot)$ depends on $p$ and is a symmetric function of the singular values of $A$.
The next theorem provides an upper bound for the variance of
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\Theta}_{2p}\coloneqq\binom{k}{p}\hat{\theta}_{2p}.}$ (5)
The bound is a function of $k$, $p$, and Schatten norms of $A$, highlighting lower variance for numerically low-rank matrices.
@@
The variance of $\hat{\Theta}_{2p}$ is bounded by
@@
$\mathrm{Var}(\hat{\theta}_{2p})\leq\frac{2p^{2}\|A\|_{4p}^{4p}}{k}+\frac{p^{2}(p-1)^{2}}{k^{2}}\left(\|A\|_{4p}^{4p}+\frac{3}{2}\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}-\frac{1}{2}\|A\|_{2p}^{4p}\right)+\mathcal{O}\left(\frac{1}{k^{3}}\right),$
where the constant in $\mathcal{O}(\cdot)$ depends on $p$ and is a symmetric function of the singular values of $A$.
Remark 5 (A comparison with (2)).
It is not straightforward to compare the variance bound of $\hat{\Theta}_{2p}$ presented in Theorem 3 with the existing bound (2), for a general matrix $A$, because different Schatten norms are involved in the bound of Theorem 3 and the ratios between different Schatten norms strongly depend on the singular value decay of the matrix. However, if the matrix $A$ is numerically low-rank, then all the Schatten norms are of comparable magnitude. If we assume, for simplicity, that all the Schatten norms are approximately $1$, we can further upper bound – and simplify – the result of Theorem 3: Using the crude bounds $c(r,\ell)\leq 3^{p}p^{p+1}$ and $\sum_{\ell=1}^{r}c(r,\ell)\leq 3^{p}p^{p+2}$, we have
$\mathrm{Var}(\hat{\theta}_{2p})$ $\lesssim\left(\frac{\binom{k}{2p}\binom{2p}{p}}{\binom{k}{p}^{2}}-1\right)+3^{p}p^{p+2}\sum_{r=1}^{p}\frac{\binom{k}{2p-r}\binom{2p-r}{p}\binom{p}{r}}{\binom{k}{p}^{2}}$
$\lesssim 0+3^{p}p^{p+2}\sum_{r=1}^{p}\frac{1}{k^{r}}\frac{p!p!}{(p-r)!r!(p-r)!}\leq 3^{p}p^{p+2}p^{2p}\frac{p}{k}=3^{p}p^{3p+3}\frac{1}{k},$ (6)
where the second inequality holds approximately, since $\binom{k}{2p-r}\binom{k}{p}^{-2}\approx k^{-r}(p!)^{2}/(2p-r)!$ when $k$ is much larger than $p$, and we totally ignore the denominator $((p-r)!)^{2}r!$ in the last line. Even with these crude approximations, the bound (6) compares favorably to (2), suggesting that the bound of Theorem 3 improves (2) at least for numerically low-rank matrices. The numerical examples (Examples 20 and 21) in Section 4 show that Theorem 3 improves (2) also in cases with no singular value decay.
Remark 6 (The constant in the $\mathcal{O}(\cdot)$).
Both in Theorem 2 and Theorem 3, the costant in the $\mathcal{O}(\cdot)$ is a function of $p$ and of the singular values of $A$. While we refrain from attempting to write down an explicit constant, we will see in the numerical examples that this constant is moderately small if the matrix $A$ has singular values with strong singular value decay, which is the focus of our paper. In this case, the first-order and second-order estimates are good approximations for the variance for moderate values of $k$. This is not true when $A$ does not exhibit singular value decay: for instance, if we consider the identity matrix (see Example 20 below), the constant in the $\mathcal{O}(\cdot)$ actually leads to imprecise first-order and second-order estimates.
Remark 7 (Rule of thumb for selecting $k$).
In the case of a numerically low-rank matrix $A$, when all Schatten norms are of comparable magnitude, Theorem 2 and Corollary 4 suggest that one needs to take $k$ to be proportional to $p^{2}$ in order to obtain a meaningful estimate for the $2p$-th power of the Schatten-$2p$ norm of $A$.
Remark 8 (A comparison with Hutchinson trace estimation).
When $p=1$, we have $\hat{\theta}_{2}=\frac{1}{k}\mathrm{trace}(\Omega^{T}A^{T}A\Omega)$, which coincides with the Hutchinson trace estimator applied to the matrix $A^{T}A$. When $p>1$, however, the two estimators differ and the variance of $\hat{\theta}_{2p}$ is usually worse than the variance of estimators based on the Hutchinson trace estimator, though both are unbiased. Specifically, since the variance of the vanilla Hutchinson trace estimator [9] for a matrix $B$ with Gaussian random vectors is $2\|B\|_{F}^{2}$, applying it to (1) gives an unbiased estimator of $\|A\|_{2p}^{2p}$ with variance
$\frac{2p}{k}\|A^{2p}\|_{F}^{2}=\frac{2p}{k}\|A\|_{4p}^{4p},$
which is smaller than (3). We remark that, in order to use the Hutchinson trace estimator with $p>1$, the matvecs with $A$ needs to be sequentially applied at least $p$ times. When the only available information about the matrix $A$ is its sketch $Y$, to the best of the authors’ knowledge, the estimator $\hat{\theta}_{2p}$ proposed in [10] is the only method to estimate the Schatten-$2p$ norms of $A$.